Fermi National Laboratory

Accelerator Update for December 5, 2007 - December 7, 2007


Wednesday December 5

The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (145.7E10) and the Antiproton Source (Pbar) stack (9.3E10), with the Main Injector (MI) and Booster operating normally, with Linac providing protons only to the TeV due to vacuum problems, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting on the next store, and with the NuMI and SciBooNE / MiniBooNE experiments waiting for the Linac vacuum problem to be resolved after the TeV shot.

At 8:47 AM, Operations prepared Linac for an access to repair the vacuum problem.

Operations established store 5780 at 9:48 AM with an initial luminosity of 122.4E30.

Operations prepared Booster for an access at 10 AM to replace weak RF station power amplifiers.

A Linac expert reported at 10:03 AM that he was shutting off the Klystron LCW distribution water skid and beginning repairs.

A Pbar expert reported at 11:24 AM that he was working on the Lithium Lens water system.

Linac experts completed their work on the Klystron water skid at 12:20 PM.

Booster came out of access at 2:01 PM.

Operations reported at 2:40 PM that various vacuum experts had checked the Linac vacuum and had not found any leaks. The vacuum pumps were turned on.

Linac experts began a beam-pulse study at 3:20 PM. The vacuum was holding at 2.4 E6 Torr.

Operations established beam to Sci/MiniBooNE at 4:54 PM.

Operations prepared Linac for an access at 6:29 PM.

At 7:02 PM, Linac experts reported that they would have to replace a vacuum pump (IPBD2).

At 9:28 PM, the Booster East Anode Power Supply tripped off on a LCW flow indication. It would not reset. Operators called in an expert.

Operations prepared Booster for access at 10:44 PM.

Vacuum experts reported at 10:56 PM that they had installed the new Linac vacuum pump (IPBD2). Operations turned the pump on at 11:22 PM.

At 11:34 PM, Operations reported that the Linac vacuum pump (IPBD2) would not stay on. Experts told operators to turn it on every half hour and run it until its temperature gets too high.

Thursday December 6

The midnight shift began with Operations monitoring store 5780 (32.61E30), the stash (3.4E10), and the stack (10.7E10), and with NuMI and Sci /MiniBooNE off due to Linac vacuum problems.

At 12:28 AM, a Linac vacuum pump (IPBD2) tripped off. Operations waited 30 minutes as directed and then turned the pump back on. This process of turning on the pump, running for about 10 minutes and then tripping off, and then waiting 30 minutes to turn on again, happened throughout the midnight shift.

Experts took over the job of operating IPBD2 at 6:32 AM.

At 10:17 AM, Operations made Linac ready for a very short access to turn off and valve off a vacuum roughing pump in the Linac dump line.

The Neutron Therapy Facility reported at 12:02 PM that they had treated their last patient for the day.

TeV experts began a crystal collimator study at 12:13 PM.

A Pbar expert began work on the Accumulator horizontal damper system at 1:17 PM.

Operations prepared MiniBooNE for an access at 1:49 PM.

While trying to solve the Linac vacuum problem at 2:52 PM, experts opened up and took a picture of the beam entrance to the Momentum dump.

LCW experts reported at 3:02 PM that they had changed the flow turbines on the Linac Klystrons.

At 3:04 PM, Electrical support experts accessed NuMI and reported that they found damaged 480-volt cables for the newly installed dehumidifiers.

Operations gave a Linac vacuum status report at 4:35 PM: the dump was isolated from the beam line and vacuum pumped. The vacuum improved, but not sufficiently enough to say the leak was in the Momentum dump. The next steps will be to check the inserted vacuum cap, the pump connection, the beam vacuum valves, and then if everything checks out okay to remove the wire scanner (DE2W).

Operations terminated store 5780 at 6:20 PM.

At 6:31 PM, Operations reported that a quadrupole magnet (C:B0Q2) quenched while it was being ramped down.

Operations prepared D0 and CDF for access at 6:46 PM.

Operations reported at 7:20 PM that experts had removed the Linac wire scanner (DE2W). Linac beam was arranged to use the Straight Ahead Dump, but with a limited amount of beam. Later, experts identified the wire scanner as the cause of the vacuum problems.

Operations established beam to Sci/MiniBooNE and Pbar at 8:33 PM.

CDF came out of access at 8:49 PM.

At 8:55 PM, Operations reported having Linac beam position problems that tripped the beam permit. This happened several times in about 40 minutes. It reset.

Operations established beam to NuMI at 9:18 PM.

D0 came out of access at 9:35 PM.

Operations put the TeV through a dry squeeze at 11:18 PM.

Operations reported at midnight that the Recycler had stashed antiprotons once during the evening shift.

Friday December 7

The midnight shift began with Operations monitoring a stash (75.7E10), with Pbar stacking (16.2E10), and with NuMI and Sci/MiniBooNE taking beam.

At 1:05 AM, KRF3 tripped off for the first of four times in about five hours. It reset.

Operations put the TeV through a dry squeeze at 2:17 AM.

Operations put the TeV through a wet squeeze at 3:24 AM.

Operations turned off NuMI beam at 4:59 AM in preparation for a scheduled access.

At 5:24 AM, KRF3 tripped off for the first of two times in nine minutes. It reset.

Operations put the TeV through a dry squeeze at 7:56 AM.

Operations reported at 7:58 AM that the Recycler had stashed four times during the midnight shift.

The Plans for Friday and the Weekend

The plan for today is to put the Linac beam line back together up to the Momentum dump without the wire scanner and then resume normal operations. This will occur around noon today after the last NTF patient treatment gets completed. The rest of the day and the weekend will consist of stacking, stashing, and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 12/07/2007    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory